OpenStudy (anonymous):

Show steps for the following equation. 25.862=-2.4(6.2+0.9v)-6.7(v-7.8)

OpenStudy (anonymous):

First, you must use the distributive law, remember also Please Excuse My Dear Aunt Sally. The first thing to do according to that is to get rid of the parenthesis, so the equation becomes: 25.862 = -14.88 - 2.16v -6.7v + 52.26 What I did to get that was the following: 25.862 = -14.88 + (-2.4*6.2) + (-2.4)*.9v + (-6.7)*v + (-6.7*-7.8) Can you solve for v from here?

OpenStudy (anonymous):

Hint: Isolate all of the terms with v in them on one side of the equation, then combine them.

OpenStudy (anonymous):

Next, combining all of the terms (anything seperated by a minus or a plus sign is a term by the way...for example in the expression x+2x+3x, x is a term, 2x is a term, and 3x is a term. So if we combine the terms with v in them in the above equation from the last step I showed you, you will end up with 25.862 + 14.88 - 52.26 = -2.16v-6.7v

OpenStudy (anonymous):

See what I did? I moved all of the terms with v in them on one side, and the terms without v on the other side. This is because now we have isolated v. Now that v is isolated on one side, we can combine the terms with v in them. This gives -11.518 =-8.86v

OpenStudy (anonymous):

Can you solve for v now?